OpenStudy (sasogeek):

the 3rd and 7th term of an arithmetic progression are -1 and 11 respectively. find the first term.

OpenStudy (anonymous):

a + 2d = -1 => 3a + 6d = -3 a + 6d = 11 solve for a

OpenStudy (ash2326):

1st....3a+6d=-3 2nd....a+6d=11 1st-2nd 2a=-14 a=-7 a+6d=11 6d=18 d=3 first term is -7 common difference 3